OpenStudy (anonymous):

Let f be a function such that f "(x) = 6x + 8. Find f(x) if the graph of f is tangent to the line 3x - y = 2 at the point (0,-2).

OpenStudy (anonymous):

good lord

OpenStudy (anonymous):

If you integrate twice, you'll get 2 constants of integration. So basically you need to find those 2 constants. f'(x)=3x^2+8x+c and f(x)=3/2 x^3 + 4x^2 + cx +k. To find c and k you need 2 equations.First the curve passes through (0,-2) and then f'(x) at (0,-2) is same as slope of the line.

OpenStudy (anonymous):

3x - y = 2 you need to do derivative

OpenStudy (anonymous):

\[f'(x)=3x^2+8x+c\] and we know at x = 0 the slope is 3 (because that is the slope of the line) so set \[f'(0)=c=3\] and we know \[c=3\] so \[f'(x)=3x^2+8x+3\]

OpenStudy (anonymous):

to find y; y;;

OpenStudy (anonymous):

now we know that \[f(x)=x^3+4x^2+3x+c\] and now we know that if x = 0 then y = -2, so solve \[f(0)=c=-2\] and therefore \[c=-2\]

OpenStudy (anonymous):

thus the function must be \[f(x)=x^2+4x^2+3x-2\]

OpenStudy (anonymous):

AWESOME...THANK YOU SO MUCH, THIS IS WHAT I GOT AS WELL